Given the thermochemical equation: 2NaN3(s) → 2 Na(s) + 3N2(g) ∆H˚= +42.7 kJ. What is the value of ∆H˚ or the preparation of 1.50 kg of N2(g)?
1
Expert's answer
2020-11-02T09:05:06-0500

According to the given equation, "\Delta{H^0}=+42.7 kJ" for 3 moles of N2 that is produced. The mass of 3 moles of N2 equals "28.01g/mol\times 3mol=84.03g."

Therefore, for 1.50 kg (or 1500 g) of N2:


"\Delta{H^0}=\frac{42.7kJ\times1500g}{84.03g}=+762kJ"


Answer: +762 kJ